# A Practical Protocol for Quantum Oblivious Transfer from One-Way Functions

### v1

@@ 1.2 Technical overview
Despite these changes, we prove that the OT functionality is simulation-based secure given an equivocal and a $(\chi,\epsilon)$-relaxed extractable bit commitment. Having a practical implementation as our final goal, we take into account the possible experimental errors as well as their correction, and we provide analytical expressions based on [BF12] that facilitates the bench-marking of the needed quantum resources. Moreover, we propose a method to distill a number of $n_{OT}$ QOT keys from a single run of the protocol. This could be done if a large number of BB84 states was required to reach a desired smaller sampling error (thus a tighter $\Delta$-security bound).
@@ 2.2 Bit commitment
Idealb: The quantum simulator outputs $(\rho^{\mathcal{C}}_{com},\rho^{\mathcal{R}}_{com})\mapsfrom\mathcal{Q}_{\mathcal{R}^{*},com}(\rho)$.
Then, $\rho^{\mathcal{R}^{*}}_{final}$mapsfrom $\langle\mathcal{Q}_{\mathcal{R}^{*},dec}(\rho^{\mathcal{C}}_{com}),\mathcal{R}_{dec}^{*}(\rho^{\mathcal{R}^{*}}_{com})\rangle$.
@@
A bit commitment protocol $(\mathcal{C},\mathcal{R})$ is statistical-binding if, for all poly-size committer $\mathcal{C}^{*}=\{\mathcal{C}_{com}^{*},\mathcal{C}_{dec}^{*}\}$ and $\lambda\in\mathbb{N}$, such that:
@@
there exists a bit $b\in\{0,1\}$ for which:
@@
Definition 4 ($(\eta,\epsilon)$-relaxed statistical binding)
@@
$\left(\boldsymbol{\rho}^{\mathcal{C}^{*}}_{com}(\lambda),\boldsymbol{\rho}^{\mathcal{R}}_{com}(\lambda)\right)=\left(\bigotimes_{i=1}^{m}\rho^{\mathcal{C}^{*}}_{com,i},\bigotimes_{i=1}^{m}\rho^{\mathcal{R}}_{com,i}\right)\mapsfrom\left(\langle\mathcal{C}^{*}_{com,i}(1^{{\lambda}},b_{i}),\mathcal{R}_{com,i}(1^{{\lambda}})\rangle\right)_{i\in m}.$
$(\mathcal{C},\mathcal{R})$ is ($\eta,\epsilon$)-relaxed-binding if there exists a set $I\subseteq[m]$, where $I\geq(1-\eta)m$, and a string $\boldsymbol{b}\in\{0,1\}^{|I|}$, if, with probability at least $1-\epsilon$, there exists a negligible function $\nu(\lambda)$ such that:
$\text{I\kern-1.49994ptP}[OUT_{\mathcal{R}}\langle\mathcal{C}_{dec}^{*}(1^{\lambda},\boldsymbol{\rho}^{\mathcal{C}^{*}}_{com}(\lambda)),\mathcal{R}_{dec}(1^{\lambda},\boldsymbol{\rho}^{\mathcal{R}}_{com}((\lambda))\rangle|_{I}\neq\boldsymbol{b}(\lambda)]\leq\nu(\lambda).$
@@
Definition 5 (($\chi$,$\epsilon$)-relaxed extractability)
Let $(\mathcal{C},\mathcal{R})$ be a bit commitment protocol. Given a sequence of auxiliary states $\left(\rho_{i},\sigma_{i}\right)_{i\in m}$ where $\rho_{i},\sigma_{i}\in\mathcal{B}_{1}(\mathbb{C}^{{\lambda}})$ and $\lambda\in\mathbb{N}$, $m$ sequential repetitions of commitments $(\rho^{\mathcal{C}^{*}}_{com,1},\rho^{\mathcal{R}}_{com,1}),$ $,(\rho^{\mathcal{C}^{*}}_{com,m},\rho^{\mathcal{R}}_{com,m})$ are ($\chi$,$\epsilon$)-relaxed-extractable if, for any poly-size quantum malicious committer $\mathcal{C}^{*}$ =$=\{(\mathcal{C}^{*}_{com,i}(1^{{\lambda}}))_{i\in m},$ $\bigotimes_{i=1}^{m}\rho_{i}\}$, there exists a quantum polynomial time extractor $\mathcal{Q}_{\mathcal{C}^{*}}=\{\mathcal{Q}_{\mathcal{C}^{*},com},\mathcal{Q}_{\mathcal{C}^{*},dec}\}$ such that, for any poly-size distinguisher $\mathcal{D}^{*}=\{(\mathcal{D}^{*}_{i})_{i\in m},\bigotimes_{i=1}^{m}\sigma_{i}\}$ and for every polynomial-size opening strategy $\left(\mathcal{C}^{*}_{dec,i}\right)_{i\in m}$, there is a negligible function $\nu(\lambda)>0$ that obeys:
$\left|\text{I\kern-1.49994ptP}[\mathcal{D}^{*}(\sigma_{\lambda},Real)=1]-\text{I\kern-1.49994ptP}[\mathcal{D}^{*}(\sigma_{\lambda},Ideal)=1]\right|=\nu(\lambda),$
@@
Real: Interaction between $\mathcal{R}_{dec}(\rho^{\mathcal{R}}_{com})$ and $\mathcal{C}^{*}_{dec}(\rho^{\mathcal{C}^{*}}_{com})$ such that $(\rho^{\mathcal{C}^{*}}_{final},b_{i})$ $\mapsfrom\langle\mathcal{C}^{*}_{dec}(\rho^{\mathcal{C}^{*}}_{com}),\mathcal{R}_{dec}(\rho^{\mathcal{R}}_{com})\rangle_{i}$, with $b_{i}\in\{0,1,\perp\}$. Then, the output after $m$ sequential repetitions is given by $\left(\boldsymbol{\rho}^{\mathcal{C}^{*}}_{final}(\lambda),\boldsymbol{b}(\lambda)\right)\mapsfrom\left(\langle\mathcal{C}^{*}_{com,i}(1^{{\lambda}},b_{i}),\mathcal{R}_{com,i}(1^{{\lambda}})\rangle\right)_{i\in m}$, where $\boldsymbol{b}\in\{0,1\}^{|I|}$ with $I\geq(1-\chi)m$.
Ideal: The simulator computes $(\boldsymbol{\rho}^{\mathcal{C}}_{com}(\lambda),\boldsymbol{\rho}^{\mathcal{R}}_{com}(\lambda),\boldsymbol{b^{*}}(\lambda))\mapsfrom(\mathcal{Q}_{\mathcal{C}^{*},com}(\rho_{i}))_{i\in\lambda}$, where $\boldsymbol{b}^{*}\in\{0,1,\mathbbold{\qm}\}^{m}$. Then, $(\boldsymbol{\rho}^{\mathcal{C}}_{final}(\lambda),\boldsymbol{b}(\lambda))$mapsfrom $\langle\mathcal{C}^{*}_{dec}(\boldsymbol{\rho}^{\mathcal{C}}_{com}(\lambda)),\mathcal{R}_{dec}(\boldsymbol{\rho}^{\mathcal{R}}_{com}(\lambda))\rangle$. The simulator outputs FAIL if :
  • |$|\{i:b^{*}_{i}=\mathbbold{\qm}\}|\geq\chi$, or
@@ 3 Equivocal Commitment
In this section, we present a compiler that yields an equivocal ($\eta$, $\epsilon$)-relaxed statistically binding bit commitment from a statistically binding computationally hiding bit commitment. As previously discussed in the Introduction 1, our compiler is directly based on [BCKM21], but we reduce the number of repetitions in order to improve efficiency at the cost of requiring relaxed binding properties. The protocol is described in Algorithm 1 and we prove the relaxed binding property in Section 3.1 and equivocality in Section 3.2.
@@ 3.1 Relaxed statistical binding
Given a statistical binding and computationally hiding commitment, we can build a commitment named EqCommitment that is ($\eta$, $\epsilon$)-relaxed statistically binding (Definition 4) and computationally hiding (Definition 1).
@@
which is equivalent to the fact that if $\text{I\kern-1.49994ptP}[E]\geq\epsilon=2^{-n\eta}$, then $|B_{0}|\leq n\eta$, which shows that EqCommitment is $(\eta,2^{-n\eta})$-relaxed statistically binding. The proof of Equation 3.1 holds since for each $i\in B_{0}$, there exists one $\gamma_{i}^{*}$ such that $u^{\gamma_{i}^{*}}_{0,i}\neq u^{\gamma_{i}^{*}}_{1,i}$, and in order for $E$ be true, the challenged $\gamma_{i}$ for that run is different than $\gamma_{i}^{*}$, which happens with probability $\frac{1}{2}$. Since all challenges are picked independently, the probability is:
@@ 3.2 Equivocality
$\mathcal{R}$ commits to $\left((\hat{x}_{i},\hat{\theta}_{i})\right)_{i\in[4\lambda_{EX}]}$ using EqCommitment.
@@
7:$\mathcal{C}$ randomly samples $2k$ disjoint subsets $M_{j}\subseteq\bar{E}$ of size $m$, with $m$ given by the size of the preimage of a 2-universal hash function $\{h_{j}(\cdot,\cdot):\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\lambda_{PQS}}\}_{j\in[2k]}$, where $k=\lfloor\lambda_{EX}/m\rfloor$.
@@
10:$\mathcal{C}$ sends $(M_{j})_{j\in[2k]}$, $\boldsymbol{\theta}$ and $(\boldsymbol{r}^{j})_{j\in[2k]}$ to $\mathcal{R}$. $\mathcal{C}$ sends the syndromes $(Synd_{j})_{j\in 2k}$ to correct a proportion $\alpha$ of errors.
@@
    • $\tilde{x}_{i}=\hat{x}_{i}$ whenever $\tilde{\theta}_{i}=\hat{\theta}_{i}$ up to an experimental error $\alpha$ and, if not, aborts.
@@
4 Equivocal and relaxed extractable commitment
We construct an equivocal and ($\chi$, $\epsilon$)-relaxed extractable commitment (Definition 5), named ERE-Commitment, based on an equivocal and ($\eta$, $\epsilon$)-relaxed statistically binding commitment. To achieve this, we first demonstrate the equivocality and relaxed statistically binding properties of the ERE-Commitment, and then proceed to prove its relaxed extractability.
4.1 ($\eta$,$2^{-k\eta}$)-relaxed statistical binding of the ERE-Commitment
Theorem 4.1
If $Com$ is a statistically binding commitment scheme, then for any function $\eta(k)$ ERE-Commitment is $\left(\eta,2^{-k\eta}\right)$-relaxed statistically binding.
Proof
The proof is similar to the one of Theorem 3.1. In ERE-Commitment, the seeds of the base commitments $Com_{r}$ are used as described in Algorithm 3. Therefore, when the honesty check is done, the sampled population is not the number of sequential repetitions of the EqCommitment $wk$ but the number of different families $k$ since the EqCommitment instances which make use of seeds derived from the same family $\boldsymbol{p}^{j}$ are opened together.
4.2 Equivocality
Theorem 4.2
Proof
The main difference with respect to the proof for EqCommitment is that the equivocal simulator has to answer correctly the commitment of $w$ equivocal boxes in parallel with single challenge bit $\gamma_{r}$. Given a computationally hiding bit commitment, $\left|p(\psi)-1/2\right|=negl(\lambda)$ also for this case. Hence, all the properties of the Algorithm 2 and the application of Watrous’ rewinding lemma are equivalent for this scheme. Therefore, from computational hiding, $Real_{b}$ and $Ideal_{b}$ of Def. 2 are indistinguishable. ∎
Corollary 1
Proof
4.3 Relaxed-extractability
  • Hybrid 0. Corresponds to the real protocol described in Algorithm 3.
@@ 4.3 Relaxed-extractability
2:For $i\in[4\lambda_{EX}]$, execute the quantum equivocal simulator $\mathcal{Q}_{\mathcal{R}^{*}}$ to generate a tuple of $((c^{\bar{\gamma}}_{\delta})^{i}_{\delta=0,1})_{i\in[4\lambda_{EX}]}$ dummy commitments. After each commitment session, $(\rho^{\mathcal{S}}_{i},\rho^{y}_{com,i})\mapsfrom\mathcal{Q}_{\mathcal{R}^{*},com}(\rho)$, where $\rho^{\mathcal{S}}_{i}$ is stored by $\mathcal{Q}_{\mathcal{C}^{*}}(\rho)$, and $\boldsymbol{\rho}^{y}_{com,i}$ is used by the malicious committer $\mathcal{C}^{*}$ as an input in the next session.
@@
4:By applying the equivocal simulator $\mathcal{Q}_{Eq}$, $\mathcal{Q}_{\mathcal{C}^{*}}$ opens the challenged bits $z_{i}\in E$ to the correct values $(\hat{x}_{i},\hat{\theta}_{i})_{i\in[E]}$. To do so, it executes sequentially for each $i\in[E]$: $\rho^{x}_{final}$mapsfrom $\langle\mathcal{Q}_{\mathcal{R}^{*},dec}(\hat{x}_{i},\hat{\theta}_{i},z_{i},\rho^{x}_{com}),\mathcal{C}_{dec}^{*}(\rho^{y}_{com})\rangle$, where $\rho^{x}_{com}$ is the current state of $\mathcal{C}^{*}$, updated after each session.
@@
Proof
Lemma 6
There exists a negligible function $\nu(\lambda)$, such that,
Proof
A quantum poly-time distinguisher $\mathcal{D}^{*}$ would be able to discriminate between both distributions if $Hyb_{2}$ and $Hyb_{3}$ had different probabilities for output $\perp$ during the honesty check. Since the honesty check is performed in the same way in both Hybrids, they are indistinguishable. ∎
Lemma 7
@@
which is equivalent to the fact that, if $\text{I\kern-1.49994ptP}[E_{1}]\geq\epsilon=2^{-\zeta k}$, then $|B_{1}|\leq\zeta k$. The proof of Equation 4.1 holds since, for each $r\in B_{1}$, there exists one set of indexes $\{\gamma^{*},\delta^{*},q^{*}\}$ such that $\boldsymbol{p}_{i_{q^{*},\delta^{*}}}^{j^{r,\gamma^{*}}}$ is not the seed of the commitment, and in order for $E_{1}$ to be true, the challenged family of commitments seeded by $\boldsymbol{p}^{j^{r,\gamma}}$ for the $r$-th set of $w$ EqCommitment instances is different from $\boldsymbol{p}^{j^{r,\gamma^{*}}}$, which happens with probability $\frac{1}{2}$. Since all challenges are picked independently, the probability that
@@
$Hyb_{3}$ and $Hyb_{4}$ would be distinguishable if the malicious committer $\mathcal{C}^{*}$ has committed to $b_{i}$ with an equivocated commitment more than a proportion $\zeta=\omega(\log^{2}(k)/k)$, or if the malicious committer $\mathcal{C}^{*}$ uses as seeds $\boldsymbol{p}^{j}\neq PRG(h(\tilde{\mathbf{x}}^{j}))$ in more than a proportion $\zeta=\omega(\log^{2}(k)/k)$ of the keys and $\text{Hyb}_{3}$ has not output $\perp$ in the honesty check phase. This happens with probability:
$\begin{split}\text{I\kern-1.49994ptP}[\text{FAIL}|\text{Ideal}]=&\text{I\kern-1.49994ptP}[\text{ number of non extractable commitments}\geq\eta+\zeta]\\ &\leq\text{I\kern-1.49994ptP}[\text{ number of non extractable commitments}\geq\zeta].\end{split}$
@@
Thus, ERE-Commitment is equivocal and ($\omega(\log^{2}(k)/k),\epsilon$)-relaxed-extractable in the sense of Definitions 2 and 5, with $\epsilon=1-negl(k)$.

### v2

@@ 1.2 Technical overview
Despite these changes, we prove that the OT functionality is simulation-based secure given an equivocal and a $\chi$-relaxed extractable bit commitment. Having a practical implementation as our final goal, we take into account the possible experimental errors as well as their correction, and we provide analytical expressions based on [BF12] that facilitates the bench-marking of the needed quantum resources. Moreover, we propose a method to distill a number of $n_{OT}$ QOT keys from a single run of the protocol. This could be done if a large number of BB84 states was required to reach a desired smaller sampling error (thus a tighter $\Delta$-security bound).
1.3 Related works
In [ABKK22], the authors also focus on reducing round complexity by leveraging the Quantum Random Oracle Model (QROM). In the QROM model, equivocality and extractability are achieved for most common constructions of commitments, so no further resources are needed. While, based on our estimations, the number of BB84 states sent in their protool is comparable  [^6] Assuming a security level of 256 bits and that the random oracle is instantiated using SHA3, the maximum number of oracle queries by the adversary would be approximately $2^{128}/(1.2\cdot 10^{6})$. Based on these parameters, we estimate that around $2\cdot 10^{6}$ BB84 states would be required in their protocol to achieve a security distance of $2^{-80}$ due to privacy amplification., this number does not account for practical errors such as as bit-flips, photon loss or multi-photon events. Therefore, we see their result and ours as complementary, since each protocol focus on minimizing a different resource. Moreover, we also highlight that our security proof is in the plain model, and it does not depend on heuristic implementations of the random oracle.
Further proposals, such as [CCLY23], consider relaxed version of simulations, where the run-time of the simulator depends polynomially on the desired security level. In our work, we can achieve negligible distinguishing probability with polynomial-time simulators.
@@ 2.2 Bit commitment
Idealb: The quantum simulator outputs $(\rho^{\mathcal{C}}_{com},\rho^{\mathcal{R}^{*}}_{com})\mapsfrom\mathcal{Q}_{\mathcal{R}^{*},com}(\rho)$.
Then, $\rho^{\mathcal{R}^{*}}_{final}$mapsfrom $\langle\mathcal{Q}_{\mathcal{R}^{*},dec}(\rho^{\mathcal{C}}_{com},b),\mathcal{R}_{dec}^{*}(\rho^{\mathcal{R}^{*}}_{com})\rangle$.
@@
A bit commitment protocol $(\mathcal{C},\mathcal{R})$ is statistical-binding if, for every unbounded-size committer $\mathcal{C}^{*}=\{\mathcal{C}_{com}^{*},\mathcal{C}_{dec}^{*}\}$ and $\lambda\in\mathbb{N}$, such that:
@@
there exists a bit $b\in\{0,1\}$ and a negligible function $\nu(\cdot)$ such that with probability at least $1-\nu(\lambda)$,
@@
Definition 4 ($\eta$-relaxed statistical binding)
@@
$\left(\boldsymbol{\rho}^{\mathcal{C}^{*}}_{com}(\lambda),\boldsymbol{\rho}^{\mathcal{R}}_{com}(\lambda)\right)\mapsfrom\left(\langle\mathcal{C}^{*}_{com,i}(1^{{\lambda}},b_{i}),\mathcal{R}_{com,i}(1^{{\lambda}})\rangle\right)_{i\in m}.$
$(\mathcal{C},\mathcal{R})$ is $\eta$-relaxed-binding if given the commitment there exists a set $I\subseteq[m]$, where $I\geq(1-\eta)m$ and a negligible function $\nu(\cdot)$ such that with probability at least $1-\nu(\lambda)$ over $\left(\boldsymbol{\rho}^{\mathcal{C}^{*}}_{com}(\lambda),\boldsymbol{\rho}^{\mathcal{R}}_{com}(\lambda)\right)$, there exists a string $\boldsymbol{b}\in\{0,1\}^{|I|}$ such that,
$\text{I\kern-1.49994ptP}[OUT_{\mathcal{R}}\langle\mathcal{C}_{dec}^{*}(1^{\lambda},\boldsymbol{\rho}^{\mathcal{C}^{*}}_{com}(\lambda)),\mathcal{R}_{dec}(1^{\lambda},\boldsymbol{\rho}^{\mathcal{R}}_{com}((\lambda))\rangle|_{I}\neq\boldsymbol{b}(\lambda)]\leq\nu(\eta).$
@@
Definition 5 ($\chi$-relaxed extractability)
Let $(\mathcal{C},\mathcal{R})$ be a bit commitment protocol. Given a sequence of auxiliary states $\left(\rho_{i},\sigma_{i}\right)_{i\in m}$ where $\rho_{i},\sigma_{i}\in\mathcal{B}_{1}(\mathbb{C}^{{\lambda}})$ and $\lambda\in\mathbb{N}$, $m$ sequential repetitions of commitments $(\rho^{\mathcal{C}^{*}}_{com,1},\rho^{\mathcal{R}}_{com,1}),$ $,(\rho^{\mathcal{C}^{*}}_{com,m},\rho^{\mathcal{R}}_{com,m})$ are $\chi$-relaxed-extractable if, for any poly-size quantum malicious committer $\mathcal{C}^{*}$ =$=\{(\mathcal{C}^{*}_{com,i}(1^{{\lambda}}))_{i\in m},$ $(\rho_{i})_{i\in m}\}$, there exists a quantum polynomial time extractor $\mathcal{Q}_{\mathcal{C}^{*}}=\{\mathcal{Q}_{\mathcal{C}^{*},com},\mathcal{Q}_{\mathcal{C}^{*},dec}\}$ such that, for any poly-size distinguisher $\mathcal{D}^{*}=\{(\mathcal{D}^{*}_{i})_{i\in m},(\sigma_{i})_{i\in m}\}$ and for every polynomial-size opening strategy $\left(\mathcal{C}^{*}_{dec,i}\right)_{i\in m}$, there is a negligible function $\nu(\chi)>0$ that obeys:
$\left|\text{I\kern-1.49994ptP}[\mathcal{D}^{*}(\sigma_{\lambda},Real)=1]-\text{I\kern-1.49994ptP}[\mathcal{D}^{*}(\sigma_{\lambda},Ideal)=1]\right|=\nu(\chi),$
@@
Real: Interaction between $\mathcal{R}_{dec}(\rho^{\mathcal{R}}_{com})$ and $\mathcal{C}^{*}_{dec}(\rho^{\mathcal{C}^{*}}_{com})$ such that $(\rho^{\mathcal{C}^{*}}_{final},b_{i})$ $\mapsfrom\langle\mathcal{C}^{*}_{dec}(\rho^{\mathcal{C}^{*}}_{com}),\mathcal{R}_{dec}(\rho^{\mathcal{R}}_{com})\rangle_{i}$, with $b_{i}\in\{0,1,\perp\}$. Then, the output after $m$ sequential repetitions is given by $\left(\boldsymbol{\rho}^{\mathcal{C}^{*}}_{final}(m),\boldsymbol{b}(m)\right)\mapsfrom\left(\langle\mathcal{C}^{*}_{com,i}(1^{{\lambda}},b_{i}),\mathcal{R}_{com,i}(1^{{\lambda}})\rangle\right)_{i\in m}$, where $\boldsymbol{b}\in\{0,1\}^{m}$.
Ideal: The simulator computes $(\boldsymbol{\rho}^{\mathcal{C}}_{com}(m),\boldsymbol{\rho}^{\mathcal{R}}_{com}(m),\boldsymbol{b^{*}}(m))\mapsfrom(\mathcal{Q}_{\mathcal{C}^{*},com}(\rho_{i}))_{i\in\lambda}$, where $\boldsymbol{b}^{*}\in\{0,1,\mathbbold{\qm}\}^{m}$. Then, $(\boldsymbol{\rho}^{\mathcal{C}}_{final}(m),\boldsymbol{b}(m))$mapsfrom $\langle\mathcal{C}^{*}_{dec}(\boldsymbol{\rho}^{\mathcal{C}}_{com}(m)),\mathcal{R}_{dec}(\boldsymbol{\rho}^{\mathcal{R}}_{com}(m))\rangle$. The simulator outputs FAIL if :
  • |$|\{i:b^{*}_{i}=\mathbbold{\qm}\}|\geq\chi m$, or
@@ 2.4 Watrous Rewinding Lemma
2.5 Error resilience
Handling error correction in our protocol involves more subtleties compared to, for example, QKD, since in our case Alice and Bob are not trusted. In quantum communication protocols, there are two primary sources of errors: noise and loss. In the case of loss, since the proposed protocol incorporates bit commitments for each BB84 state, the loss of a state will result in the absence of a commitment for that instance of the protocol. Discarding non-committed instances has no impact on the security analysis of the protocol.
The potential presence of malicious parties prevents them from cooperating to estimate the noise. Therefore, an error correction subroutine must be included in the protocol to handle errors without any collaboration between parties [DFSS06, ENG+14, LPAK23]. An $(\mathcal{M},K,t)$-error correction code can correct errors in a code $C$ composed of $K$ elements from a metric space $\mathcal{M}$, such that for every element $w\in\mathcal{M}$, there exists at most one codeword $c\in C$ within a ball of radius $t$ around $w$.
We propose the use of a non-interactive (classical) error correction code, which is syndrome-based, such as Low-Density Parity-Check (LDPC) codes [Gal60]. The syndrome of an element $w$ is its projection onto the orthogonal subspace of the code, capturing all the necessary information for decoding.
Finally, we consider the possibility of multiphoton events. While an ideal BB84 source generates single photons, experimental sources could accidentally generate more than one copy of a BB84 instance. We take into account this leakage of information for the privacy amplification of our proposed protocol. We present this as a general variable $\vartheta$ that accounts for the information leakage.
@@ 3 Equivocal Commitment
In this section, we present a compiler that yields an equivocal $\eta$-relaxed statistically binding bit commitment from a statistically binding computationally hiding bit commitment. As previously discussed in the Introduction, our compiler is directly based on [BCKM21], but we reduce the number of repetitions in order to improve efficiency at the cost of requiring relaxed binding properties. The protocol is described in Algorithm 1 and we prove the relaxed binding property in Section 3.1 and equivocality in Section 3.2.
@@ 3.1 Relaxed statistical binding
Given a statistical binding and computationally hiding commitment, we can build a commitment named EqCommitment that is $\eta$-relaxed statistically binding (Definition 4) and computationally hiding (Definition 1).
@@
which is equivalent to the fact that if $\text{I\kern-1.49994ptP}[E]\geq 2^{-n\eta}$, then $|B_{0}|\leq n\eta$, which shows that EqCommitment is $(\eta,2^{-n\eta})$-relaxed statistically binding. The proof of Equation 3.1 holds since for each $i\in B_{0}$, there exists one $\gamma_{i}^{*}$ such that $u^{\gamma_{i}^{*}}_{0,i}\neq u^{\gamma_{i}^{*}}_{1,i}$, and in order for $E$ be true, the challenged $\gamma_{i}$ for that run is different than $\gamma_{i}^{*}$, which happens with probability $\frac{1}{2}$. Since all challenges are picked independently, the probability is:
@@ 3.2 Equivocality
4 Equivocal and relaxed extractable commitment
We construct an equivocal and $\chi$-relaxed extractable commitment (Definition 5), named ERE-Commitment, based on an equivocal and $\eta$-relaxed statistically binding commitment.
We present our protocol in Algorithm 3. We first notice that if the committer and receiver are honest, then the protocol succeeds.
Lemma 4
Given an implementation of ERE-Commmitment for which the noise error in the BB84 transmission is $\delta\leq\alpha$ between an honest committer and an honest receiver, the committer does not abort.
Proof
As discussed in Section 2.5, an $(\mathcal{M},4\lambda_{EX},4\alpha\lambda_{EX})$-error correction code can unambiguously correct up to a proportion of $\alpha$ errors. Therefore, for any $\delta\leq\alpha$, there is at most one codeword inside the radius of the error correction code, not aborting in Step 11.3 of Algorithm 3
We now prove the equivocality and the relaxed extractability of our protocol.
4.1 Equivocality
Theorem 4.1
Proof
The main difference with respect to the proof for EqCommitment is that the equivocal simulator has to answer correctly the commitment of $w$ equivocal boxes in parallel with single challenge bit $\gamma_{r}$. Given a computationally hiding bit commitment, $\left|p(\psi)-1/2\right|=negl(\lambda)$ also for this case. Hence, all the properties of the Algorithm 2 and the application of Watrous’ rewinding lemma are equivalent for this scheme. Therefore, from computational hiding, $Real_{b}$ and $Ideal_{b}$ of Def. 2 are indistinguishable.
Corollary 1
Proof
4.2 Relaxed-extractability
Before proving relaxed-extractability, we first show that our commitment scheme is relaxed-binding.
Theorem 4.2
If $Com$ is a statistically binding commitment scheme, then for any function $\eta(k)$ ERE-Commitment is $\eta$-relaxed statistically binding.
Proof
The proof is similar to the one of Theorem 3.1. In ERE-Commitment, the seeds of the base commitments $Com_{r}$ are used as described in Algorithm 3. Therefore, when the honesty check is done, the sampled population is not the number of sequential repetitions of the EqCommitment $wk$ but the number of different families $k$ since the EqCommitment instances which make use of seeds derived from the same family $\boldsymbol{p}^{j}$ are opened together. ∎
  • Hybrid 0. Corresponds to the real protocol described in Algorithm 3.
@@ 4.2 Relaxed-extractability
$\mathcal{R}$ sequentially commits to $\left((\hat{x}_{i},\hat{\theta}_{i})\right)_{i\in[4\lambda_{EX}]}$ using EqCommitment.
@@
7:$\mathcal{C}$ divides $\bar{E}$ into $2k$ consecutive disjoint subsets $M_{j}\subseteq\bar{E}$ of size $m$, with $m$ given by the size of the preimage of a 2-universal hash function $\{h_{j}(\cdot,\cdot):\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\lambda_{PQS}}\}_{j\in[2k]}$, where $k=\lfloor\lambda_{EX}/m\rfloor$.
@@
10:$\mathcal{C}$ sends $(M_{j})_{j\in[2k]}$, $\boldsymbol{\theta}$ and $(\boldsymbol{r}^{j})_{j\in[2k]}$ to $\mathcal{R}$. $\mathcal{C}$ sends the syndromes $(Synd_{j})_{j\in[2k]}$ to correct a proportion $\alpha$ of errors.
@@
    • $\tilde{x}_{i}=\tilde{\hat{x}}_{i}$ whenever $\tilde{\theta}_{i}=\hat{\theta}_{i}$, where $\tilde{\hat{x}}_{i}$ are $\mathcal{R}^{\prime}s$ measures corrected with the syndromes $(Synd_{j})_{j\in[2k]}$ and, if not, aborts.
@@
Lemma 6
Proof
Lemma 7
There exists a negligible function $\nu(\lambda)$, such that,
Proof
A quantum poly-time distinguisher $\mathcal{D}^{*}$ would be able to discriminate between both distributions if $Hyb_{2}$ and $Hyb_{3}$ had different probabilities for output $\perp$ during the honesty check. Since the honesty check is performed in the same way in both Hybrids, taking into account in the error correction, they are indistinguishable. ∎
@@
2:For $i\in[4\lambda_{EX}]$, execute sequentially the quantum equivocal simulator $\mathcal{Q}_{\mathcal{R}^{*}}$ to generate a tuple of $((c^{\bar{\gamma}}_{\delta})^{i}_{\delta=0,1})_{i\in[4\lambda_{EX}]}$ dummy commitments. After each commitment session, $(\rho^{\mathcal{S}}_{i},\rho^{y}_{com,i})\mapsfrom\mathcal{Q}_{\mathcal{R}^{*},com}(\rho)$, where $\rho^{\mathcal{S}}_{i}$ is stored by $\mathcal{Q}_{\mathcal{C}^{*}}(\rho)$, and $\boldsymbol{\rho}^{y}_{com,i}$ is used by the malicious committer $\mathcal{C}^{*}$ as an input in the next session.
@@
4:By applying the equivocal simulator $\mathcal{Q}_{Eq}$, $\mathcal{Q}_{\mathcal{C}^{*}}$ opens the challenged bits $z_{i}\in E$ to the correct values $(\hat{x}_{i},\hat{\theta}_{i})_{i\in[E]}$. To do so, it executes for each $i\in[E]$: $\rho^{x}_{final}$mapsfrom $\langle\mathcal{Q}_{\mathcal{R}^{*},dec}(\hat{x}_{i},\hat{\theta}_{i},z_{i},\rho^{x}_{com}),\mathcal{C}_{dec}^{*}(\rho^{y}_{com})\rangle$, where $\rho^{x}_{com}$ is the current state of $\mathcal{C}^{*}$, updated after each session.
@@
which is equivalent to the fact that, if $\text{I\kern-1.49994ptP}[E_{1}]\geq 2^{-\zeta k}$, then $|B_{1}|\leq\zeta k$. The proof of Equation 4.1 holds since, for each $r\in B_{1}$, there exists one set of indexes $\{\gamma^{*},\delta^{*},q^{*}\}$ such that $\boldsymbol{p}_{i_{q^{*},\delta^{*}}}^{j^{r,\gamma^{*}}}$ is not the seed of the commitment, and in order for $E_{1}$ to be true, the challenged family of commitments seeded by $\boldsymbol{p}^{j^{r,\gamma}}$ for the $r$-th set of $w$ EqCommitment instances is different from $\boldsymbol{p}^{j^{r,\gamma^{*}}}$, which happens with probability $\frac{1}{2}$. Since all challenges are picked independently, the probability that
@@
$Hyb_{3}$ and $Hyb_{4}$ would be distinguishable if the malicious committer $\mathcal{C}^{*}$ has committed to $b_{i}$ with an equivocated commitment more than a proportion $\eta=\omega(\log^{2}(k)/k)$, or if the malicious committer $\mathcal{C}^{*}$ uses as seeds $\boldsymbol{p}^{j}\neq PRG(h(\tilde{\mathbf{x}}^{j}))$ in more than a proportion $\zeta=\omega(\log^{2}(k)/k)$ of the keys and $\text{Hyb}_{3}$ has not output $\perp$ in the honesty check phase. This happens with probability:
$\begin{split}\text{I\kern-1.49994ptP}[\text{FAIL}|\text{Ideal}]=&\text{I\kern-1.49994ptP}[\text{proportion of non extractable commitments}\geq\eta+\zeta]\\ &\leq\text{I\kern-1.49994ptP}[\text{ proportion of non extractable commitments}\geq\zeta].\end{split}$
@@
Thus, ERE-Commitment is equivocal and $\omega(\log^{2}(k)/k)$-relaxed-extractable in the sense of Definitions 2 and 5.